### v2

# Topotactical Hydrogen Induced Single-Band -Wave Superconductivity in La2NiO4

Abstract
La2NiO4 is an antiferromagnetic insulator with a structural resemblance to its cuprate counterpart, La2CuO4. However, La2CuO4 has a Cu2+ or 3 electronic configuration that needs to be hole or electron doped for superconductivity, whereas La2NiO4 is 3 with divalent Ni2+. Making a cuprate analog through conventional electron doping is impractical due to the rarity of tetravalent substituents for trivalent La. Here, we propose an alternative route: intercalating topotactical hydrogen, which is possible through electric-field-controlled protonation and transforms La2NiO4 into a 3 single-band two-dimensional antiferromagnetic Mott insulator analogous to La2CuO4. This we find through density-functional theory and dynamical mean-field theory calculations. The furthergoing dynamical vertex approximation predicts that H-La2NiO4 can host -wave superconductivity under 15% hole doping with a critical temperature above 20 K. Our findings not only suggest a new method for tuning the electronic structure of layered nickelates but also provide theoretical evidence for a new nickelate superconductor, awaiting experimental synthesis.
Introduction—The discovery of new superconductors Norman (2016) is one of the central goals in condensed matter physics, as each breakthrough not only deepens our fundamental understanding of the pairing mechanism but also paves the way toward achieving room-temperature superconductivity. Among the various families of superconductors, the cuprates Bednorz and Müller (1986); Anderson (1997); Tsuei and Kirtley (2000); Damascelli et al. (2003)—characterized by a single-band Fermi surface originating from - hybridization—hold the record for unconventional superconductivity at ambient pressure with critical temperatures exceeding 130 K Schilling et al. (1993). This remarkable performance has motivated extensive efforts to identify new superconducting systems that share similar electronic characteristics with cuprates.
In recent years, the quest for cupratelike superconductivity has extended to nickelates Allan et al. (1989); Bates and Eldridge (1989); Acrivos et al. (1994); Lee and Pickett (2004); Anisimov et al. (1999); Chaloupka and Khaliullin (2008); Hansmann et al. (2009); Guo et al. (2018); Zhou et al. (2018). Early attempts in this direction were unsuccessful until 2019, when Li and Hwang et al. realized superconductivity in Sr-doped NdNiO2 via a reduction process from (Nd,Sr)NiO3 by CaH2 Li et al. (2019a, 2020a). This breakthrough subsequently led to the discovery of a range of superconducting nickelates, including infinite-layer (La,Ca)NiO2 Zeng et al. (2022), (La,Sr)NiO2 Osada et al. (2021), (Pr,Sr)NiO2 Osada et al. (2020), (Nd,Ca)NiO2, finite-layer Nd6Ni5O12 Pan et al. (2022); Yan et al. (2025), pressurized (La,Pr)3Ni2O7 Sun et al. (2023); Hou et al. (2023); Wang et al. (2024a), pressurized (La,Pr)4Ni3O10 Zhu et al. (2024); Li et al. (2024); Pei et al. (2024); Zhang et al. (2025); Huang et al. (2024), and La3Ni2O7 under ambient pressure and compressive strain Ko et al. (2025); Zhou et al. (2024); Li et al. (2025a). These developments have restarted the search for nickelate and even palladate superconductors Kitatani et al. (2023) that mirror the single-band physics observed in cuprates.
Although various theoretical models have been proposed for both infinite-layer Kitatani et al. (2020); Lechermann (2020a, b); Zhang et al. (2020a); Werner and Hoshino (2020); Zhang et al. (2020b); Sakakibara et al. (2020); Zhang and Vishwanath (2020); Nomura et al. (2019); Zhou et al. (2020); Xie et al. (2022); Li and Louie (2024) and finite-layer nickelate Worm et al. (2022a); Lechermann (2022); LaBollita and Botana (2022); LaBollita et al. (2022); Carrasco Álvarez et al. (2024); Craco et al. (2022) superconductors—ranging from single-band to multiband descriptions—recent THz experiments Cheng et al. (2024, 2025) and resonant inelastic x-ray scattering (RIXS) measurements Worm et al. (2024) indicate that infinite-layer nickelate superconductors exhibit an electronic structure with a doped Hubbard single band of character similar to that of two-dimensional cuprate superconductors, suggesting -wave superconductivity. For the pressurized (La,Pr)3Ni2O7 Sun et al. (2023); Hou et al. (2023); Wang et al. (2024a) and pressurized (La,Pr)4Ni3O10 Zhu et al. (2024); Li et al. (2024); Pei et al. (2024); Zhang et al. (2025); Huang et al. (2024) systems, the incorporation of an apical oxygen atom leads to Ni-3 orbital contributions near the Fermi surface. Theoretical studies suggest that the superconductivity in these pressured samples may result either from a Zhang-Rice-like - hybridization between (in-plane) O-/ and Ni-3 Sun et al. (2023), or -wave pairing Zhang et al. (2024a); Yang et al. (2024); Zhang et al. (2023, 2024b); Yang et al. (2023); Zhang et al. (2024c); Yang et al. (2024); Liu et al. (2023); Zhang et al. (2024c) induced by a charge or spin density wave Ni et al. (2025); LaBollita et al. (2024); Li et al. (2025b); Chen et al. (2024); Samarakoon et al. (2023); LaBollita et al. (2024); Yi et al. (2024); Meng et al. (2024); Kumar et al. (2020); Liu et al. (2024); Leonov (2024).
Among various nickelates, La2NiO4 has long attracted interest due to its structural similarity to La2CuO4 Allan et al. (1989); Acrivos et al. (1994); Lander et al. (1989); Pintschovius et al. (1989); Brown (1992); Rodriguez-Carvajal et al. (1991); Wang et al. (2024b); Biało et al. (2024); Cui et al. (2024), the parent compound of the cuprate superconductors. While the undoped La2NiO4 is an antiferromagnetic Mott insulator exhibiting intriguing quantum states Guo and Temmerman (1988); Rodriguez-Carvajal et al. (1991); Lander et al. (1989); Mahadevan et al. (1997), including charge and orbital ordering Homes et al. (2007); Hücker et al. (2004); Jestädt et al. (1999); Jung et al. (2001); Mahadevan et al. (1997)—especially under hole doping—the Ni ions in La2NiO4 stay in a Ni2+ state or 3 electronic configuration. Achieving superconductivity akin to that in cuprates requires an effective 3 ( 0.2) electronic configuration Anderson (1997); Tsuei and Kirtley (2000); Scalapino (1995), which, in principle, demands excessive electron doping. However, electron doping in La2NiO4 is impractical due to the absence of tetravalent elements capable of substituting La.
An alternative and promising route is proton evolution which is realized by applying ionic liquid gating and which recently has proven effective in controlling the electronic and magnetic properties of correlated oxides such as SrRuO3 Li et al. (2020b), SrCoO2.5 Li et al. (2019b); Lu et al. (2022), CaRuO3 Shen et al. (2021), and Ca2RuO4 Tian et al. (2024) that are structurally analogous to La2NiO4. In these systems, the intercalated proton acts as an electron donor by forming (OH)-. Motivated by these advances of liquid gating, we investigate the effect of proton evolution in La2NiO4.
To this end, we employ a combination of density-functional theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965), dynamical mean-field theory (DMFT) Metzner and Vollhardt (1989); Georges et al. (1996); Kotliar and Vollhardt (2004); Held (2007), and dynamical vertex approximation (DA) Held et al. (2008); Galler et al. (2019); Rohringer et al. (2018) as diagrammatic extensions of DMFT. We show that out-of-plane intercalated hydrogen (H) in La2NiO4 forms a proton that transfers an electron to Ni, converting Ni2+ (3) to Ni+ (3). This creates a single-band structure analogous to that in La2CuO4. Our DA calculations further predict that this protonated phase, H-La2NiO4, is in close proximity to the infinite-layer nickelate regime and is promising to exhibit superconductivity with a critical temperature above 20 K. These findings suggest that proton evolution provides a viable path for engineering superconductivity in nickelate systems and opens new avenues for exploring unconventional superconductivity beyond the cuprate family.
Method—DFT-level structural relaxations and magnetic and electronic band structure calculations were performed using the Vasp Kresse and Furthmüller (1996, 1996) and Wien2K Blaha et al. (2001); Schwarz et al. (2002) codes with the Perdew-Burke-Ernzerhof version of the generalized gradient approximation (GGA-PBE) Perdew et al. (1996) and a dense 13135 -mesh for the unit cell of La2NiO4 with and without intercalated H. The interaction parameters for single-band Ni-3 and full Ni-3 orbitals were obtained from the constrained random phase approximation (cRPA) Miyake and Aryasetiawan (2008) implemented in VASP Si et al. (2025). Ab initio molecular dynamics (AIMD) simulations implemented in VASP Kresse and Hafner (1993a, 1994, b) were performed to study the finite-temperature structural and dynamic properties of hydrided La2NiO4 (H-La2NiO4). The AIMD simulations were conducted in the canonical ensemble using a Nos-Hoover thermostat (NVT) to maintain constant temperature. A time step of 1 fs was used to accurately capture the atomic motions, and each simulation was run for a total of 103 fs to ensure adequate statistical sampling of the phase space. A large 221 supercell of H-La2NiO4 was employed in AIMD to minimize finite-size effects.
Starting points for the DMFT calculations are the Wien2K bands of single Ni-3 and full Ni-3 orbitals around Fermi level projected onto the maximally localized Wannier functions Wannier (1937); Marzari et al. (2012) using the Wien2Wannier Mostofi et al. (2008); Kuneš et al. (2010) interface. The thus obtained DFT-Wannier Hamiltonian is supplemented by a local Kanamori interaction, and we employ the fully localized limit as the double counting correction Liechtenstein et al. (1995); Anisimov et al. (1993). The resulting many-body Hamiltonian is solved at room temperature (300 K) within DMFT employing the continuous-time quantum Monte Carlo solver in the hybridization expansions Gull et al. (2011) of the W2dynamics code Parragh et al. (2012); Wallerberger et al. (2019).
Real-frequency spectra are obtained via analytic continuation using the maximum entropy method (MaxEnt) Gubernatis et al. (1991); Sandvik (1998) as implemented in the ana_cont code Kaufmann and Held (2023).
The superconducting is calculated with the DA Galler et al. (2019); Held et al. (2008); Rohringer et al. (2018). First, we compute the particle-particle vertex incorporating spin fluctuations from both the particle-hole and transversal particle-hole channels. Next, we extract the leading superconducting eigenvalues in the particle-particle channel Kitatani et al. (2019). This procedure effectively represents the first iteration of the particle-particle channel in a full parquet DA calculation. For further technical details, see Kitatani et al. (2022).
Structure—We first investigate the doping position of intercalated H in the Ruddlesden-Popper (RP) phase of La2NiO4. The undoped La2NiO4 (space group , No. 139) is illustrated in Fig. 1(a). Figures 1(b) and 1(c) present the relaxed structures of H-La2NiO4 with two favorable sites for H intercalation: out-of-plane (OOP) [Fig. 1(b)] and in-plane (IP) [Fig. 1(c)] position. First-principles DFT structural optimizations reveal that the out-of-plane configuration has a total energy of –111.98 eV per unit cell [Fig. 1(b)] compared to –111.42 eV [Fig. 1(c)] for the in-plane configuration, indicating that the out-of-plane H is more stable by 280 meV per H atom 111In the present study, we initially designed several possible structures of H-La2NiO4 by intercalating hydrogen into La2NiO4, guided by crystal symmetry analysis. Subsequent nonmagnetic DFT calculations were performed to optimize these structures and determine the ground state of H-La2NiO4. However, a more comprehensive structural search may require exploring a broader range of possible configurations within larger crystal cells. Our primary focus in this research is to investigate the electron doping effect introduced by hydrogen intercalation in La2NiO4. For details of DFT calculations for other possible structures of H-La2NiO4, see Supplemental Material (SM) SM Sec. I.
| Materials | (Å) | (Å) | Magnetism | (experimental) | (theory) | |||||
| La2CuO4 (this work) | 3.814 | 13.145 | -0.444 | -0.055 | 0.081 | 2.68 | 6.0 | -AFM | 30 K Bednorz and Müller (1986) | — |
| (Nd,La)NiO2 Kitatani et al. (2020); Si et al. (2020) | 3.890 | 3.337 | -0.395 | -0.240 | 0.118 | 2.60 | 6.5 | -AFM | 20 K Lee et al. (2023) | 36 K Kitatani et al. (2020) |
| H-La2NiO4 (this work) | 3.913 | 13.712 | -0.356 | -0.234 | 0.105 | 2.81 | 7.9 | -AFM | — | 20 K |
Since DFT calculations are performed at zero temperature, we validated the stability at finite temperature using AIMD (embedded in VASP Kresse and Hafner (1993a, 1994, b)). Figure 2 in SM SM Sec. II shows the crystal at 300 K. Notably, DFT optimization and MD simulations at 300 K yield H-O bond lengths of 0.99 Å in DFT and 0.99–1.05 Å in MD at 300 K SM , respectively, consistent with typical H-O bond lengths. Thus, the out-of-plane position doped H forms a stable, negatively charged hydroxide (OH)- with connection to the out-of-plane O. In previous work on infinite-layer nickelate LaNiO2 with H Tran and Blaha (2009); Si et al. (2023); Han et al. (2023), DFT and DMFT calculations revealed that H ions form a -type bond with Ni via the Ni-3 and H-1 orbital, resulting in an H- valence state and the formation of a Ni2+ (3) configuration [see Fig. 1(d)]. For out-of-plane H-La2NiO4 [Fig. 1(b)], which has lower energy in DFT, an opening question is whether electron doping can be achieved by intercalating H.
DFT bands—To address this question, we first performed non-spin-polarized DFT calculations for the RP-phase La2CuO4—the parent compound of single-band cuprate superconductor—and La2NiO4, a two-band AFM Mott insulator. Figure 2(a) shows that La2CuO4 features a single-band Fermi surface. In contrast, La2NiO4 exhibits a two-band Fermi surface, with Ni-3 and 3 orbitals contributing a holelike Fermi surface and an electron-type pocket, respectively. In H-La2NiO4, however, the Fermi surface displays a single-band feature as the band shifts down compared to its position in La2NiO4 [Fig. 2(b) and 2 (c)]. This indicates that the H ion adopts an H+ valence state. More specifically, in H-La2NiO4, H forms hydroxide (OH)- with O [Fig. 1(d)], resulting in a 3 (Ni+) configuration and consequently a characteristic single-band Fermi surface [Fig. 2(d)], remarkably similar to that of the cuprate La2CuO4 and infinite-layer LaNiO2. SM SM Secs. III–V demonstrate that this single-Ni-band picture is robust against both the position of topotactic H and strain. Remarkably, the latter can lead to a self-doping akin to infinite-layer nickelates.
Additionally, we calculated the electron hopping parameters of H-La2NiO4 using the Wannier projection method Marzari and Vanderbilt (1997); Souza et al. (2001); Marzari et al. (2012), including the first (), second (), and third () nearest-neighbor hoppings, as well as the ratios / and /. Our findings indicate that these hopping parameters in H-La2NiO4 closely resemble those found in infinite-layer (La,Nd)NiO2 (Table 1), such as / –0.24. Previous studies Kitatani et al. (2023) have demonstrated that these hopping parameters play a crucial role in determining the superconducting properties within the two-dimensional Hubbard model framework, especially for nickelates and palladates. They influence not only the emergence of superconductivity but also the maximum superconducting transition temperature () achievable in these materials. The similarity in hopping parameters suggests that H-La2NiO4 may exhibit comparable superconducting characteristics to (La,Nd)NiO2. Moreover, our DFT+ magnetic order calculations indicate that the ground state of H-La2NiO4 is an antiferromagnetic Mott insulator (see SM SM Sec. VI). This finding aligns with the behavior observed (experimentally) in parent compounds of cuprate superconductors, such as La2CuO4, further highlighting the parallels between these material systems.
As a consequence, both DFT band calculations and Wannier projections reveal notable electronic structure similarities between H-La2NiO4, LaNiO2 and La2CuO4. These materials exhibit features consistent with the one-band Hubbard model, which has been pivotal in understanding high-temperature superconductivity. This resemblance again underscores the potential of H-La2NiO4 as a candidate for high-temperature superconductivity, akin to its nickelate counterparts.
Correlations—DFT calculations often underestimate electronic correlations, particularly the dynamical correlation effects within 3 orbitals. Therefore, we performed DMFT calculations to examine the electronic structure near the Fermi level in greater detail. VASP-cRPA calculations for H-La2NiO4 were performed using a multiband and single-band model, respectively: for the full- model, we obtain an intraorbital =2.92 eV, Hund’s =0.62 eV, and interorbital =-2=1.68 eV; for the single-band model the Hubbard is =2.8 eV, leading to /7.9 (: the nearest intraorbital hoppings of this single orbital are summarized in Table 1).
Figure 3(a) compares the -integrated DMFT spectral functions () for multi- and single-orbital calculations. In the multiorbital (), the and orbitals are fully occupied, while the orbital is half-filled and exhibits significant renormalization, with an effective mass of approximately 3.3. This value is close to that reported for infinite-layer nickelates (4.4) Kitatani et al. (2020) and for some cuprates, it also ranges from 3 to 4 Uemura et al. (1989); Padilla et al. (2005); Worm et al. (2022b). Notably, the single-band () shown in Fig. 3(a) exhibits remarkable similarity to the full- multibands (), which strongly supports the validity of a minimal low-energy model based solely on the orbital in H-La2NiO4. Additionally, we computed the -resolved spectral function (,) for the orbital [Fig. 3(b)], revealing that the Fermi surface is entirely derived from this band, with the upper and lower Hubbard bands appearing at approximately 2 eV and –2 eV, respectively. Notably, at and points, the Ni-3 band exhibits nearly identical energies, indicating its two-dimensional (2D) character. This arises from the substantial spatial separation of Ni ions along the axis, which minimizes interlayer interactions/hopping and reinforces its 2D character. DFT+ calculations for a 221 supercell of H-La2NiO4 indicate the ground state is a -type AFM Mott-insulating state (see SM SM Sec. VI).
Estimating —The above DFT and DMFT results demonstrate the effectiveness of a minimal single-band model and its essentially 2D character. Similar to the Hubbard model for cuprates and nickelates, this single band captures the low-energy electronic structure and prospectively the superconducting properties. To further investigate superconductivity, more sophisticated many-body methods than DMFT, such as DA method Rohringer et al. (2018), must be employed. DA starts from a local irreducible vertex (depending on three frequencies and obtained at DMFT convergence) and builds nonlocal vertices and self-energies through ladder or parquet diagrams. The starting vertex in DA is nonperturbative, unlike in RPA or FLEX. This approach incorporates both the local DMFT correlations and the nonlocal spin, charge, and superconducting fluctuations, enabling the calculation of quantum critical exponents, pseudogaps, and superconducting properties in the one-band Hubbard model. Previous applications Kitatani et al. (2019, 2020, 2023); Worm et al. (2022a, 2024) have proven its power in predicting superconducting materials materials and estimating critical temperatures in cuprates and nickelates.
Figure 4(a) displays the DA Fermi surface at various hole dopings (10%, 15% and 20%) and temperatures (46 K and 116 K), based on the constructed 2D Hubbard model with the hopping parameters from Table 1. The Fermi surface is obtained from the lowest Matsubara frequency. Please note that the interlayer (-direction) hopping between orbital is justified by the spacing layers, which makes H-La2NiO4 an even more ideal two-dimensional system than infinite-layer Nd(La)NiO2 (which has a hopping of 34 meV compared to 2 meV in H-La2NiO4). For a doping level =0.85 (corresponding to H-La1.85Sr0.15NiO4), a well-defined holelike Fermi surface can be observed in Fig. 4(a). In contrast, at =0.9 (H-La1.9Sr0.1NiO4), strong AFM spin fluctuations induce significant scattering at the antinodal point =(,0). This signals the vicinity of the pseudogap regime. At the lower temperature of 46 K, the bands become more well-defined across the Brillouin zone, indicating a suppression of the hole contributions to, e.g., Hall coefficient at higher temperatures.
The superconducting critical temperature can be determined from the divergence of the superconducting susceptibility , or equivalently, the leading superconducting eigenvalue . In matrix notation Kitatani et al. (2019), =/[1-], where is the bare superconducting susceptibility and is the irreducible vertex in the particle-particle channel calculated by DA. When the leading eigenvalue of approaches 1, the superconducting susceptibility diverges. Figure 4(b) plots versus temperature at different fillings of the Ni- orbital. It shows that approaches 1 at low temperatures for =0.80 and =0.85, whereas for =0.90, there seems to be no superconductivity. Specifically, the superconducting state is absent for =0.9 down to the lowest measured temperature, and for =0.8, it is only observed at T13 K. This suggests a superconducting dome within the range 0.90n0.80. The momentum structure of the leading eigenvector (not shown) indicates -wave superconductivity.
Conclusion—The recent discovery of infinite-layer nickelate superconductors has revealed a wide range of crystal structures, chemical compositions, and potential superconducting mechanisms in different nickelates. In contrast, although La2NiO4 has been extensively studied and exhibits a variety of novel quantum states, superconductivity has not been found yet experimentally. Here, we investigated the structural, thermal, magnetic, electronic, and superconducting properties of H-La2NiO4 where hydrogen is topotactically intercalated, using state-of-the-art theoretical methods. These methods include DA, which is well-suited for calculating superconducting temperatures in two-dimensional strongly correlated oxides such as cuprates and nickelates. Unlike in LaNiO2H, it acts as an electronic donor in H-La2NiO4. This leads to a Ni+ valence with an electronic configuration similar to Cu2+. Also the in-plane AFM Mott insulating state for full hydrogen intercalation resembles that of cuprate superconductors such as La2CuO4. With an optimal hole doping of approximately 15%, our simulations predict that H-La2NiO4 exhibits a -wave superconducting state with up to 20 K, akin to its infinite-layer nickelate analogs. Against the background of recent experimental progress to add hydrogen via ionic liquid gating in perovskite and RP-oxides, the experimental synthesis of the suggested H-La2NiO4 appears possible and promising.
Acknowledgments—We thank Prof. Pu Yu, Viktor Christiansson and Eric Jacob for helpful insightful discussions. L. S. acknowledges support from the National Natural Science Foundation of China (Grant No. 12422407). L. S. and K. H. acknowledge funding through the Austrian Science Funds (FWF) Project Grant DOI 10.55776/I5398. W.-F. W. acknowledges the funding from the China Scholarship Council (CSC). Calculations have been done on the the National Supercomputing Center (Xi’an) in Northwest University and the Vienna Scientific Cluster (VSC).